# Is Your Paper Being Reviewed by an LLM? Benchmarking AI Text Detection in Peer Review

### v2

@@ 1 Introduction
Recent advancements in large language models (LLMs) have enabled their application to a broad range of domains, where LLMs have demonstrated the ability to produce plausible and authoritative responses to queries even in highly technical subject areas. These advancements have coincided with a surge in interest in AI research, resulting in increased paper submissions to leading AI conferences [1]. Consequently, workloads for peer reviewers have also increased significantly, which could make LLMs an appealing tool for lessening the burden of fulfilling their peer review obligations [2, 3, 4].
@@
Despite their impressive capabilities, the use of LLMs in the peer review process raises several ethical and methodological concerns which could compromise the integrity of the publication process [5, 6, 7, 8]. Reviewers are selected based on their expertise in a technical domain related to a submitted manuscript, which is necessary to critically evaluate the proposed research. Offloading this responsibility to an LLM circumvents the role that reviewer selection plays in ensuring proper vetting of a manuscript. Furthermore, LLMs are prone to hallucination and may not possess the ability to rigorously evaluate research publications. Therefore, the use of LLMs in an undisclosed manner in peer review poses a significant ethical concern that could undermine confidence in this important process.
Motivating the need for evaluation resources and detection tools to address this problem is the apparent increase in AI-generated text among peer reviews submitted to recent AI research conferences. Prior studies revealed an upward trend in AI-generated texts among peer reviews [9, 6]. This trend is particularly concerning given that evaluations from human and AI reviewers are not aligned [10, 6, 11] and that LLM-generated reviews lack robustness [12], suggesting the unregulated and undisclosed use of LLMs in peer review could undermine the integrity of the current system.
Despite the growing recognition of this problem, there is a lack of existing dataset resources for comprehensively evaluating the performance of AI text detection methods in the domain of peer review. To address this deficiency, we introduce the largest dataset to-date of parallel human-written and LLM-written peer reviews for 8 years of papers submitted to two leading AI research conferences, NeurIPS and ICLR (Figure 1). Our dataset consolidates human-written peer reviews from existing sources with AI-written peer reviews that we generated for the same paper using five state-of-the-art LLMs: GPT-4o [13], Claude Sonnet 3.5 [14], Gemini 1.5 pro [15], Qwen 2.5 72b [16], and Llama 3.1 70b [17]. In total, our dataset contains 788,984 peer reviews, evenly balanced between human-written reviews and AI-generated peer reviews created by these five LLMs.
We use our dataset to investigate the suitability of various AI text detection methods for identifying LLM generations in the peer review process. While limited prior work has analyzed the presence of AI-generated text in peer reviews at the corpus level [9] or has analyzed the use of propriety solutions [6], our study is the first to investigate the detectability LLM generations at the individual review level using synthetically generated AI samples, which is necessary to address this problem in practice. Specifically, we evaluate 18 existing open-source methods for AI text detection.
@@ 2 Related Work
Several datasets have been introduced to evaluate AI text detection models. RAID-TD  [18] provides a large-scale benchmark designed to assess text detection under adversarial conditions, ensuring robustness against manipulated AI-generated content. The M4 Dataset  [19] expands the scope by incorporating reviews from multiple LLMs across different languages, offering a more diverse linguistic evaluation. The HC3 Dataset  [20] consists of responses from ChatGPT and human experts, covering specialized domains such as finance, medicine, and law, in addition to general open-domain content. In contrast, the GPT Reddit Dataset (GRiD)  [21] focuses on social media conversations, compiling a diverse set of human- and AI-generated responses to Reddit discussions. Meanwhile, Beemo [22] introduces a benchmark of expert-edited machine-generated outputs, spanning creative writing, summarization, and other practical applications. These benchmarks primarily evaluate AI-generated text from a single model and do not address the domain of AI text in peer review. In contrast, our dataset is larger than most existing datasets (788k generations) and is unique in its focus on AI text detection in peer review.
@@
AI-generated text detection has been framed as a binary classification task to distinguish human-written from machine-generated text  [23, 24, 25, 26]. Solaiman et al. [27] used a bag-of-words model with logistic regression for GPT-2 detection, while fine-tuned language models like RoBERTa [28] improved accuracy  [29, 30, 31]. Zero-shot methods based on perplexity and entropy emerged as alternatives  [32, 31]. Other studies focused on linguistic patterns and syntactic features for model-agnostic detection  [30, 31]. Watermarking techniques, such as DetectGPT  [26], have also been proposed for proactive identification. Centralized frameworks like MGTBench  [33] and its refined version, IMGTB  [34], provide standardized evaluations for AI text detection. IMGTB categorizes methods into model-based and metric-based approaches. Model-based methods leverage large language models such as ChatGPT-turbo  [35] and Claude  [14]. Metric-based methods, including Log-Likelihood  [27], Rank  [31], Entropy  [31], DetectGPT  [26], and LRR  [36], rely on log-likelihood and ranking for classification.
@@
Recent studies have explored the role of LLMs in peer review, examining their influence on reviewing practices  [9, 37], simulating multi-turn interactions  [38], and assessing their reviewing capabilities  [8]. Tools like OpenReviewer  [39] provide AI-assisted review improvements, while other works focus on LLM transparency  [2] and distinguishing AI-generated content  [40]. Recent studies have investigated AI-driven review systems  [41], agentic frameworks  [42], and comparative analyses of LLM-generated reviews  [43], along with broader explorations of LLMs’ roles and limitations in peer review  [44, 45, 46, 37]. While it is not the primary focus of our work, we analyze the quality of LLM-generated peer reviews in Section 5.
@@ 3.1 Human reviews
We used the OpenReview API [47] to collect submitted manuscripts and their reviews for the ICLR conferences from 2019 to 2024, as well as for NeurIPS conferences from 2021 to 2024111NeurIPS 2020 reviews are not publicly available. Additionally, we used the ASAP dataset [48] to collect manuscripts and reviews for ICLR 2017 to 2018 and NeurIPS 2016 to 2019.
@@ 3.3 Dataset Statistics
Table 1: Dataset statistics. Each subset in the dataset is balanced, containing an equal number of human-written and AI-generated peer reviews. The calibration and test sets include reviews generated by five different large language models (LLMs), whereas the extended set includes reviews generated by GPT-4o and Llama 3.1.
@@ 4.1 Fully AI-Written Review Detectability
Prior work has shown that AUC is not necessarily reflective of how models perform at very low FPR values [49, 50, 51]. Therefore, we also report the actual TPR and FPR achieved by different detection methods at discrete low values of target FPR (0.1%, 0.5%, and 1%), which we believe to be of greatest interest for practical applications. The target FPR is used to calculate each method’s classification threshold using our calibration dataset, with the actual TPR and FPR computed over the withheld test dataset. To simulate a more challenging evaluation setting where some of the test reviews are “out-of-domain” in the sense that they come from a different conference than the calibration dataset, we use only the ICLR reviews to calibrate each method (see Section C.2 for in-domain evaluations).
@@
Table 2 provides these results separately for the detection of GPT-4o, Gemini, and Claude reviews (Llama and Qwen results are provided in Appendix C.3). Other baseline methods that failed achieve a TPR of at least 1% at a target FPR of 1% (i.e., TPR $<1\%$ at target FPR $=1\%$) are omitted.
The results in Table 2 show that among the 18 baselines, Binoculars generally achieves the best performance overall. For example, its TPR reaches 45–85% at a 1% target FPR, while other models reach only up to 19% TPR. However, even with the best-performing Binoculars, performance drops significantly at more stringent FPR levels—for instance, it achieves only 17% TPR at 0.1% FPR for GPT-4o reviews. These results highlight a key challenge in AI-generated peer review detection: the difficulty of robustly identifying AI-written text at the level of individual peer reviews. At the same time, they point to the need for new methods that can improve accuracy and robustness under low-FPR constraints. To this end, we explore a purpose-built detection method (Anchor) in Section 5.1, which leverages additional context specific to the peer review detection task.
@@ 5 Analysis
5.1 Can manuscript context improve detection of AI-generated peer reviews?
To test this idea, we introduce a method which utilizes the manuscript by comparing the semantic similarity between a test reviews (TR) and a synthetic "Anchor Review" (AR) generated for the same manuscript. The AR can be generated by any LLM. We use a simple, generic prompt (Appendix E.2) to generate the AR without prior knowledge of the user prompts (i.e., the AR prompt differs from those used to create reviews in the testing dataset). Once an AR is generated for a given paper (Eq.1) and a testing review (TR) is provided, we obtain their embeddings using a text embedding model (EM, Eqs.2 and 3). The semantic similarity between the embeddings of the AR and TR is then computed using the cosine similarity function (Eq.4). Finally, this similarity score is compared against a learned threshold ($\theta$): if the score exceeds the threshold, the review is classified as AI-generated; otherwise, it is not (Eq. 5). Putting everything together, the method is formalized as:
$\text{AR}=\text{LLM}(\text{paper},\text{Prompt}_{\text{AR}})$ (1)
[…]
Compared to the 18 baseline models, our anchor embedding approach consistently achieves the highest AUC for AI peer review (Figure 2) and outperforms other baseline methods overall (Table 2). The performance gap is especially notable for GPT-4o, which is generally harder to detect; for example, at a target FPR of 0.1%, the anchor embedding method achieves a TPR of 63.5% for GPT-4o reviews, compared to 17.1% for the next-best method (Binoculars), which is an absolute improvement of 46.4%. This indicates that the anchor embedding approach is particularly effective for the most challenging review samples. On Gemini reviews, where baseline methods already perform well, the anchor embedding method performs similarly to Binoculars. These results suggest that leveraging the context of the submitted manuscript through anchor embeddings offers substantial benefits for high-risk, low-FPR detection tasks, especially when reviews are written by the most advanced LLMs.
@@ 5.2 Is AI peer review detection robust to prompt variations?
We further investigated prompt-induced variation using sentence-level embedding analysis. Figure S2 visualizes a t-SNE projection of review embeddings from two groups: (i) AI-generated reviews using the primary score-aligned prompt and (ii) AI-generated reviews using the archetype prompt. No meaningful separation emerged between these two review types, suggesting they share broadly similar linguistic and semantic characteristics. This finding reinforces the robustness of our prompting strategy and implies that moderate prompt variation does not lead to drastic stylistic shifts in model outputs.
@@ 5.3 Can detection models distinguish human reviews edited by LLMs?
LLMs are widely used for writing assistance tasks such as grammar correction and fluency enhancement [52, 53], especially benefiting non-native speakers. However, excessive reliance on LLMs can lead to substantial rewriting, blurring the line between human- and AI-authored reviews. To simulate this scenario, we took human-written peer reviews from the ICLR 2021 GPT-4o test subset and generated AI-edited versions at four increasing levels of editing: Minimum, Moderate, Extensive, and Maximum (see Appendix E.3 for prompts). To validate that the edits are semantically distinct, we measured cosine similarity between the original and edited reviews using a sentence embedding model. As shown in Table S12, similarity scores decrease as the level of editing increases.
We evaluate detection models on two fronts: (i) their ability to flag edited reviews as AI-generated, and (ii) their ability to rank the reviews by degree of AI involvement using Normalized Discounted Cumulative Gain (NDCG) [54]. Table 3 compares our anchor embedding method with Binoculars, the strongest-performing baseline from previous experiments (as shown in Table 2). Both models achieve high NDCG scores (0.90 for Anchor, 0.86 for Binoculars), indicating that they can generally rank reviews according to the degree of AI editing. However, their flagging behavior differs. At the 0.1% FPR threshold, both models rarely flag minimally to extensively edited reviews as AI-generated ($\leq$2.5%), but diverge sharply at the Maximum level: Anchor flags 60.8% of these reviews, compared to just 9.2% for Binoculars. This trend persists at the 1% FPR threshold, with generally higher flagging rates across all levels. These results suggest that the Anchor method is more sensitive to high levels of AI involvement, whereas Binoculars is more conservative at the upper end.
Overall, these results suggest that high-performing detection models can distinguish varying levels of AI editing in human-written reviews. However, lightly edited texts remain difficult to detect, highlighting a challenge for future work on identifying hybrid human–AI content.

### v3

@@ 1 Introduction
Recent advancements in large language models (LLMs) have enabled their application to a broad range of domains, where LLMs have demonstrated the ability to produce plausible and authoritative responses to queries even in highly technical subject areas. These advancements have coincided with a surge in interest in AI research, resulting in increased paper submissions to leading AI conferences (Audibert et al., 2022). Consequently, workloads for peer reviewers have also increased significantly, which could make LLMs an appealing tool for lessening the burden of fulfilling their peer review obligations (Kuznetsov et al., 2024; Kousha & Thelwall, 2024; Zhuang et al., 2025).
@@
Despite their impressive capabilities, the use of LLMs in the peer review process raises several ethical and methodological concerns which could compromise the integrity of the publication process (Hosseini & Horbach, 2023; Latona et al., 2024; Seghier, 2024; Zhou et al., 2024). Reviewers are selected based on their expertise in a technical domain related to a submitted manuscript, which is necessary to critically evaluate the proposed research. Offloading this responsibility to an LLM circumvents the role that reviewer selection plays in ensuring proper vetting of a manuscript. Furthermore, LLMs are prone to hallucination and may not possess the ability to rigorously evaluate research publications. Therefore, the use of LLMs in an undisclosed manner in peer review poses a significant ethical concern that could undermine confidence in this important process.
Motivating the need for evaluation resources and detection tools to address this problem is the apparent increase in AI-generated text among peer reviews submitted to recent AI research conferences. Prior studies revealed an upward trend in AI-generated texts among peer reviews (Liang et al., 2024a; Latona et al., 2024). This trend is particularly concerning given that evaluations from human and AI reviewers are not aligned (Drori & Te’eni, 2024; Latona et al., 2024; Ye et al., 2024) and that LLM-generated reviews lack robustness (Li et al., 2025), suggesting the unregulated and undisclosed use of LLMs in peer review could undermine the integrity of the current system.
Despite the growing recognition of this problem, there is a lack of existing dataset resources for comprehensively evaluating the performance of AI text detection methods in the domain of peer review. To address this deficiency, we introduce the largest dataset to-date of parallel human-written and LLM-written peer reviews for 8 years of papers submitted to two leading AI research conferences, NeurIPS and ICLR (Figure 1). Our dataset consolidates human-written peer reviews from existing sources with AI-written peer reviews that we generated for the same paper using five state-of-the-art LLMs: GPT-4o Achiam et al. (2023), Claude Sonnet 3.5 Anthropic (2023), Gemini 1.5 pro Team et al. (2023), Qwen 2.5 72b Bai et al. (2023), and Llama 3.1 70b Dubey et al. (2024). In total, our dataset contains 788,984 peer reviews, evenly balanced between human-written reviews and AI-generated peer reviews created by these five LLMs.
We use our dataset to investigate the suitability of various AI text detection methods for identifying LLM generations in the peer review process. While limited prior work has analyzed the presence of AI-generated text in peer reviews at the corpus level (Liang et al., 2024a) or has analyzed the use of propriety solutions (Latona et al., 2024), our study is the first to investigate the detectability LLM generations at the individual review level using synthetically generated AI samples, which is necessary to address this problem in practice. Specifically, we evaluate 18 existing open-source methods for AI text detection.
@@ 2 Related Work
Several datasets have been introduced to evaluate AI text detection models. RAID-TD  (Dugan et al., 2024) provides a large-scale benchmark designed to assess text detection under adversarial conditions, ensuring robustness against manipulated AI-generated content. The M4 Dataset  (Wang et al., 2024) expands the scope by incorporating reviews from multiple LLMs across different languages, offering a more diverse linguistic evaluation. The HC3 Dataset  (Guo et al., 2023) consists of responses from ChatGPT and human experts, covering specialized domains such as finance, medicine, and law, in addition to general open-domain content. In contrast, the GPT Reddit Dataset (GRiD)  (Qazi et al., 2024) focuses on social media conversations, compiling a diverse set of human- and AI-generated responses to Reddit discussions. Meanwhile, Beemo (Artemova et al., 2024) introduces a benchmark of expert-edited machine-generated outputs, spanning creative writing, summarization, and other practical applications. These benchmarks primarily evaluate AI-generated text from a single model and do not address the domain of AI text in peer review. In contrast, our dataset is larger than most existing datasets (788k generations) and is unique in its focus on AI text detection in peer review.
@@
AI-generated text detection has been framed as a binary classification task to distinguish human-written from machine-generated text  (Bakhtin et al., 2019; Jawahar et al., 2020; Fagni et al., 2021; Mitchell et al., 2023b). Solaiman et al. (2019) used a bag-of-words model with logistic regression for GPT-2 detection, while fine-tuned language models like RoBERTa (Liu et al., 1907) improved accuracy  (Zellers et al., 2019; Uchendu et al., 2020; Gehrmann et al., 2019). Zero-shot methods based on perplexity and entropy emerged as alternatives  (Ippolito et al., 2020; Gehrmann et al., 2019). Other studies focused on linguistic patterns and syntactic features for model-agnostic detection  (Uchendu et al., 2020; Gehrmann et al., 2019). Watermarking techniques, such as DetectGPT  (Mitchell et al., 2023b), have also been proposed for proactive identification. Centralized frameworks like MGTBench  (He et al., 2023) and its refined version, IMGTB  (Spiegel & Macko, 2023), provide standardized evaluations for AI text detection. IMGTB categorizes methods into model-based and metric-based approaches. Model-based methods leverage large language models such as ChatGPT-turbo  (OpenAI, 2023) and Claude  (Anthropic, 2023). Metric-based methods, including Log-Likelihood  (Solaiman et al., 2019), Rank  (Gehrmann et al., 2019), Entropy  (Gehrmann et al., 2019), DetectGPT  (Mitchell et al., 2023b), and DetectLLM  (Su et al., ), rely on log-likelihood and ranking for classification.
While there is some similarity between our Anchor approach and methods such as DetectLLM and DNA-GPT (Yang et al., 2023), there are key differences in both the functionality and applicability of our method relative to this prior work. DetectLLM assumes a white-box detection setting where the detector has access to the LLM used to generate the evaluated text. Unfortunately this setting is unrealistic for our task of detecting AI text in peer review because there is no way of knowing which LLM was used to generate the review. Additionally, we view frontier commercial models as the most likely to be used in practice due to their superior capabilities; since these models are only available via an API, methods such as DetectLLM which require full access to the model are not compatible.
Another key difference between our Anchor method and both DetectLLM and DNA-GPT is that it leverages additional context which is available in the peer review setting: the manuscript of the paper which is being reviewed. In contrast, DNA-GPT truncates the middle of an evaluated text, regenerates the remainder with an LLM, and then analyzes the difference between the original & regenerated portions via n-gram analysis (in the black-box detection setting). This is unlikely to produce text which is as representative of AI-generated reviews because it lacks grounding in the source content (the paper). To the best of our knowledge, our anchor embedding approach is unique among AI text detection methods in its use of the additional context available in the peer review setting.
@@
Recent studies have explored the role of LLMs in peer review, examining their influence on reviewing practices  (Liang et al., 2024a; b), simulating multi-turn interactions  (Tan et al., 2024), and assessing their reviewing capabilities  (Zhou et al., 2024). Tools like OpenReviewer  (Tyser et al., ) provide AI-assisted review improvements, while other works focus on LLM transparency  (Kuznetsov et al., 2024) and distinguishing AI-generated content  (Mosca et al., 2023). Recent studies have investigated AI-driven review systems  (Tyser et al., 2024), agentic frameworks  (D’Arcy et al., 2024), and comparative analyses of LLM-generated reviews  (Liu & Shah, 2023), along with broader explorations of LLMs’ roles and limitations in peer review  (Jin et al., 2024; Santu et al., 2024; Sukpanichnant et al., 2024; Liang et al., 2024b). While it is not the primary focus of our work, we analyze the quality of LLM-generated peer reviews in Section 5.
@@ 3.1 Human reviews
We used the OpenReview API (OpenReview, ) to collect submitted manuscripts and their reviews for the ICLR conferences from 2019 to 2024, as well as for NeurIPS conferences from 2021 to 2024111NeurIPS 2020 reviews are not publicly available. Additionally, we used the ASAP dataset Yuan et al. (2022) to collect manuscripts and reviews for ICLR 2017 to 2018 and NeurIPS 2016 to 2019.
@@ 3.2 AI reviews
Throughout the course of our study, we evaluated multiple prompting strategies ranging from prompts which produce fully AI-written reviews to ones which request varying levels of LLM editing of human-written peer reviews (Section 5.2). Our primary focus is on the former scenario of detecting fully AI-written peer review, which we believe to be the most pressing ethical problem. However, we also included the LLM-edited human reviews (with varying levels of editing) from our analysis in Section 5.2 in our released dataset.
@@ 3.3 Dataset Statistics
Table 1: Dataset statistics. Each subset in the dataset is balanced, containing an equal number of human-written and AI-generated peer reviews. The calibration and test sets include reviews generated by five LLMs, whereas the extended set includes reviews generated by GPT-4o and Llama 3.1.
@@ 4 Experimental Results
4.1 Improving detection of AI-generated peer reviews using manuscript context
To test this idea, we introduce a method which utilizes the manuscript by comparing the semantic similarity between a test reviews (TR) and a synthetic "Anchor Review" (AR) generated for the same manuscript. The AR can be generated by any LLM. We use a simple, generic prompt (Appendix E.2) to generate the AR without prior knowledge of the user prompts (i.e., the AR prompt differs from those used to create reviews in the testing dataset). Once an AR is generated for a given paper (Eq.1) and a testing review (TR) is provided, we obtain their embeddings using a text embedding model (EM, Eqs.2 and 3). The semantic similarity between the embeddings of the AR and TR is then computed via cosine similarity (Eq.4). Finally, this similarity score is compared against a learned threshold ($\theta$): if the score exceeds the threshold, the review is classified as AI-generated (Eq. 5).
(Formalization of Anchor Method)
$\text{AR}=\text{LLM}(\text{paper},\text{Prompt}_{\text{AR}})$ (1)
$\text{Emb}_{\text{AR}}=\text{EM(AR)}$ (2)
$\text{Emb}_{\text{TR}}=\text{EM(TR)}$ (3)
$\text{Score}=\text{Cosine\_similarity}(\text{Emb}_{\text{AR}},\text{Emb}_{\text{TR}})$ (4)
@@ 4.2 Fully AI-Written Review Detectability
Prior work has shown that AUC is not necessarily reflective of how models perform at very low FPR values (Yang et al., 2023; Krishna et al., 2024; Tufts et al., ). Therefore, we also report the actual TPR and FPR achieved by different detection methods at discrete low values of target FPR (0.1%, 0.5%, and 1%), which we believe to be of greatest interest for practical applications. The target FPR is used to calculate each method’s classification threshold using our calibration dataset, with the actual TPR and FPR computed over the withheld test dataset. To simulate a more challenging evaluation setting where some of the test reviews are “out-of-domain” in the sense that they come from a different conference than the calibration dataset, we use only the ICLR reviews to calibrate each method (see Section C.2 for in-domain evaluations).
@@
Table 2 provides these results separately for the detection of GPT-4o, Gemini, and Claude reviews (Llama and Qwen results are provided in Appendix C.3). Other baseline methods that failed achieve a TPR of at least 1% at a target FPR of 1% (i.e., TPR $1\%$ at target FPR $1\%$) are omitted.
The results in Table 2 show that among the 18 baselines, Binoculars generally achieves the best performance overall. For example, its TPR reaches 45–85% at a 1% target FPR, while other models reach only up to 19% TPR. However, even with the best-performing Binoculars, performance drops significantly at more stringent FPR levels—for instance, it achieves only 17% TPR at 0.1% FPR for GPT-4o reviews. These results highlight a key challenge in AI-generated peer review detection: the difficulty of robustly identifying AI-written text at the level of individual peer reviews. At the same time, they point to the need for new methods that can improve accuracy and robustness under low-FPR constraints. To this end, we also evaluate a purpose-built detection method (Anchor) as described previously in Section 4.1, which leverages additional context specific to the peer review detection task.
Compared to the 18 baseline models, our Anchor approach consistently achieves the highest AUC for AI peer review (Figure 2) and outperforms other baseline methods overall (Table 2). The performance gap is especially notable for GPT-4o, which is generally harder to detect; for example, at a target FPR of 0.1%, the anchor embedding method achieves a TPR of 63.5% for GPT-4o reviews, compared to 17.1% for the next-best method (Binoculars), which is an absolute improvement of 46.4%. This indicates that the anchor embedding approach is particularly effective for the most challenging review samples. On Gemini reviews, where baseline methods already perform well, the anchor embedding method performs similarly to Binoculars. These results suggest that leveraging the context of the submitted manuscript through anchor embeddings offers substantial benefits for high-risk, low-FPR detection tasks, especially when reviews are written by the most advanced LLMs.
@@ 5.1 Is AI peer review detection robust to prompt variations?
We also investigated prompt-induced variation using sentence-level embedding analysis. Figure S2 visualizes a t-SNE projection of review embeddings from two groups: (i) AI-generated reviews using the primary score-aligned prompt and (ii) AI-generated reviews using the archetype prompt. No meaningful separation emerged between these two review types, suggesting they share broadly similar linguistic and semantic characteristics. This finding reinforces the robustness of our prompting strategy and implies that moderate prompt variation does not lead to drastic stylistic shifts in model outputs.
Evaluation on Agent-Based and Structured Generation Pipelines
To assess whether more complex prompting strategies fundamentally alter the detectability of AI-generated peer reviews, we evaluated Binoculars on reviews generated by several recent agent-based or structured LLM review systems: AI Scientist (Lu et al., 2024), AgentReview (Jin et al., 2024), and DeepReview (Zhu et al., 2025). These systems incorporate features like multi-step workflows, reviewer personas, and self-reflection. We paired these AI-generated reviews with human-written reviews from ICLR 2022 and measured detectability using AUROC (full results in Appendix D.3). Despite the increased sophistication of these generation pipelines, we found Binoculars consistently achieve high AUROC (>> 0.99), which is $\sim$0.03 larger than its performance on our own GPT-4o-generated baseline. This suggests that such advanced prompting or agentic strategies do not fundamentally degrade detectability, supporting the broader applicability of our benchmark.
@@ 5.2 Can detection models distinguish human reviews edited by LLMs?
LLMs are widely used for writing assistance tasks such as grammar correction and fluency enhancement Laban et al. (2024); Raheja et al. (2024), especially benefiting non-native speakers. However, excessive reliance on LLMs can lead to substantial rewriting, blurring the line between human- and AI-authored reviews. To simulate this scenario, we took human-written peer reviews from the ICLR 2021 GPT-4o test subset and generated AI-edited versions at four increasing levels of editing: Minimum, Moderate, Extensive, and Maximum (see Appendix E.3 for prompts). To validate that the edits are semantically distinct, we measured cosine similarity between the original and edited reviews using a sentence embedding model. As shown in Table S14, similarity scores decrease as the level of editing increases.
We evaluate detection models on two fronts: (i) their ability to flag edited reviews as AI-generated, and (ii) their ability to rank the reviews by degree of AI involvement using Normalized Discounted Cumulative Gain (NDCG) Järvelin & Kekäläinen (2002). Table 3 compares our anchor embedding method with Binoculars, the strongest-performing baseline from previous experiments (as shown in Table 2). Both models achieve high NDCG scores (0.90 for Anchor, 0.86 for Binoculars), indicating that they can generally rank reviews according to the degree of AI editing. However, their flagging behavior differs. At the 0.1% FPR threshold, both models rarely flag minimally to extensively edited reviews as AI-generated ($\leq$2.5%), but diverge sharply at the Maximum level: Anchor flags 60.8% of these reviews, compared to just 9.2% for Binoculars. This trend persists at the 1% FPR threshold, with generally higher flagging rates across all levels. These results suggest that the Anchor method is more sensitive to high levels of AI involvement, whereas Binoculars is more conservative at the upper end.
Overall, these results suggest that high-performing detection models can distinguish varying levels of AI editing in human-written reviews. However, lightly edited texts remain difficult to detect, highlighting a challenge for future work on identifying hybrid human–AI content. Additionally, the inverse of this analysis (i.e., “human revises AI draft”) could be a promising direction to investigate in future work; we expect to see an inverse relationship in this setting as our "AI edits human draft" results, where greater levels of human editing will reduce the likelihood of the review being flagged as AI-written.
@@
5.3 Detectability of mixed-authorship peer reviews
In addition to investigating the detectability of human-written and LLM-edited peer reviews, we also conducted an analysis of a different mixed-authorship scenario where AI peer reviews are generated from bullet points (see Lee et al. (2024) for related approaches). Specifically, we aimed to simulate the scenario of a human drafting bullet points and using an LLM to assist with writing the full review. We did this by first asking an LLM (GPT-4.1) to condense a human-written peer review into a series of bullet points which capture the core points of the review using the following prompt: Given a peer review for an academic research paper, condense the core points of the review into bullet points. Output only the bullet points with no additional commentary or formatting. We then prompted GPT-4.1 to generate a review solely from the resulting bullet points (which themselves were derived from human-written reviews sampled from the ICLR 2021 test set) using the following prompt: Given a set of bullet points describing a peer review of an academic research paper, write the full text of the peer review.
Table 4: Anchor detectability results for AI reviews generated from bullet points
Target FPR Actual FPR Actual TPR
0.1% 0.2% 5.8%
0.5% 0.8% 14.4%
1.0% 2.2% 18.6%
The Anchor embedding method detection results for the 500 AI reviews generated from bullet points are provided in Table 4. Consistent with our previous analysis of mixed-authorship generation settings (AI editing of human-written peer reviews), we find that these reviews which are generated from the core points of human-written peer reviews are much more challenging to detect. This generation setting could be viewed as a legitimate use case for LLMs in the peer review process because the core points and ideas of the review are still originating from a human, with the LLM serving as a writing assistant. Our main focus is the identification of unethical uses of LLMs in the peer review process (i.e., when it is used to fully generate the peer review); these results demonstrate how our Anchor method performs well at detecting such uses while having relatively low detection rates for cases where LLMs are primarily used for writing assistance.